find the area of the triangle formed by the line with the equation 2x - 5y + 20 = 0 and the x and y axes.

Answer by Alan3354(69443)

find the area of the triangle formed by the line with the equation 2x - 5y + 20 = 0 and the x and y axes.
----------------
Find the x & y intercepts.
2x - 5y + 20 = 0
(0,4) & (-10,0)
It's a right triangle with legs of 4 and 10
Area = b*h/2 = 20 sq units